OpenStudy (anonymous):

Can someone check the next post I will make in this question and tell me if my answers are correct

OpenStudy (anonymous):

Change into logarithm form: 0.01=10^-2 I change it into log_10^0.01=-2

OpenStudy (anonymous):

Change into exponential form: log_7^191.322=2.7 I changed it to 7^2.7=191.322

OpenStudy (loser66):

yup for both